S8-3 Use target costing to analyze data (Learning Objective 2) See the Winter Sports Inc. data from S8-2. Assume that Winter Sports’ reputation has diminished and other resorts in the vicinity are charging only $65 per lift ticket. Winter Sports has become a price-taker and won’t be able to charge more than its competitors. At the market price, Winter Sports’ managers believe they will still serve 750,000 skiers and snowboarders each season. If Winter Sports can’t reduce its costs, what profit will it earn? State your answer in dollars and as a percent of assets. Will investors be happy with the profit level? Show your analysis. Assume that Winter Sports has found ways to cut its fixed costs to $30 million. What is its new target variable cost per skier/snowboarder? Assume investors want to earn a 15% return on assets. Compare this to the current variable cost per skier/ snowboarder. Comment on your results

data from 8-2

S8-2 Determine pricing approach and target price (Learning Objective 2) Winter Sports Inc. operates a Rocky Mountain ski resort. The company is planning its lift ticket pricing for the coming ski season. Investors would like to earn a 15% return on the company’s $100 million of assets. The company incurs primarily fixed costs to groom the runs and operate the lifts. Winter Sports projects fixed costs to be $33,750,000 for the ski season. The resort serves 750,000 skiers and snowboarders each season. Variable costs are $10 per guest. Currently, the resort has such a favorable reputation among skiers and snowboarders that it has some control over the lift ticket prices."
Answer #1

S8_2.

Let the target price be p.

Revenue. 750000p

Fixed cost. (33750000)

Variable cost. 750000x10. (7500000)

______________________

Profit. 750000p - 41250000

Therefore    750000p - 41250000 = 100x106x0.15

or p = $75

S8_3

.

Revenue. 750000x65 48750000

Fixed cost. (33750000)

Variable cost. 750000x10. (7500000)

______________

$7500000 I.e. ( 7500000 / 100x106) = 0.075 or 7.5 % of assets

If Winter Sports can reduce its fixed cost to $30 million then the profit would be

48750000 - 30000000 - Variable cost

or 18750000 - variable cost

This should be equal to the targeted return of 15 % on assets

therefore   18750000 - variable cost = 100x106x0.15

\Rightarrow target variable cost = 3750000

or targeted cost per guest = 3750000 / 7500000

=$5 this is lower than the current cost of $10 per guest

Thus Winter Sports will have to find ways to reduce its varable cost per guest to half in order to acheive the targeted return on the assets.